OpenStudy (anonymous):

(20x^4+10x^3-49x^2+11) divided by (5x^2-1)

OpenStudy (tkhunny):

The first term in the result is 4x^2. What's next?

OpenStudy (anonymous):

is it supposed to be 5x^2

OpenStudy (anonymous):

I don't get it. ._.

OpenStudy (anonymous):

oh well what im talking about is if its supposed to be divided by (5x^2 -1)

OpenStudy (tkhunny):

\(\dfrac{20x^{4}}{5x^{2}} = 4x^{2}\) The first term in the result is \(4x^2\). \(\dfrac{10x^{3}}{5x^{2}} = 2x\) The second term in the result is \(2x\). The third term is a little trickier.

OpenStudy (zehanz):

You have to do a long division: First write it like this: (on the dots the answer will gradually appear) 5x²-1 / 20x^4 +10x³ - 49x² + 11 \ ... Then ask yourself what number times 5x² will be 20x^4. (we ignore the "-1" part of 5x²-1. It will be allright) That number is 4x², because 5x² times 4x² = 20x^4. So write 4x² on the dots and multiply, including the "-1". Put the result below 20x^4+10x^3 and subtract: 5x²-1 / 20x^4 +10x³ - 49x² + 11 \ 4x² 20x^4 - 4x² - ---------------------- 10x³ - 45x² + 11 Now do the same trick again: 5x² times 2x is 10x³: 5x²-1 / 20x^4 +10x³ - 49x² + 11 \ 4x² + 2x 20x^4 - 4x² - ---------------------- 10x³ - 45x² + 11 10x² -2x - ---------------- -45x² +2x +11 And again: 5x² times -9 is -45x²: 5x²-1 / 20x^4 +10x³ - 49x² + 11 \ 4x² + 2x - 9 20x^4 - 4x² - ---------------------- 10x³ - 45x² + 11 10x² -2x - ---------------- -45x² +2x +11 -45x² +9 - -------------- 2x +2 We're done! The remainder is 2x+2, the quotient is 4x²+2x+9.